# Generalized Euler Equation from Effective Action: Implications for the Smarr Formula in AdS Black Holes

### v2

@@ 5 Comments on cosmological constant as a thermodynamic variable
where $d$ is the dimension of the bulk spacetime [^9] For reviews on extended black hole thermodynamics, see Mann4 ; Mann5 . Other relevant literature on the subject includes Mann2 ; Mann3 ; Clifford2 ; Kastor1 .. Clearly, the notion of pressure $p_{\Lambda}$ does not correspond to the pressure of the dual field theory. For this reason, the physical interpretation of the Smarr formula was investigated from a holographic perspective in Karch .
@@ 6 Discussion
It’s well-known that the Schwarzian action captures the near-horizon dynamics of near-extremal black holes Amhed ; Stanford . In contrast, our effective action given in (40) is an off-shell formulation that captures not only the one-point functions (the stress tensor), but also the Ward identities (the Navier-Stokes equations without dissipation). In this way, following closely the work of Me , it should be possible to derive a decoupling limit from the action (67), which encodes the dynamics of a holographic charged fluid, and connect it to the one-dimensional anomalous perfect fluid of Jensen2 . In fact, the author of Me already achieved this decoupling limit for a restricted class of CFT3, although using a different set of variables from the fluid variables.
@@
Similarly, the isometries of the degenerate target metric $G_{MN}$ given in (24) are closely related to Carrollian diffeomorphisms. Motivated by this, it should be possible to formulate an effective action for Carrollian fluids CR1 ; CR2 ; CR3 ; CR4 ; CR5 ; CR6 ; CR7 ; CR8 . In this context, Carrollian diffeomorphisms would serve both as gauge symmetries of the effective action and as spacetime symmetries of the fluid’s background [^10] This perspective aligns naturally with the Gauge/Gravity duality Liu5 , as it likely follows from taking a flat limit of the fluid. There is also a body of work exploring the connection between Carrollian fluids and horizon symmetries CR9 ; CR10 ; CR11 ; CR12 ; CR13 .. With this action principle in place, the Smarr formula for asymptotically flat black holes should follow in a manner similar to the derivation presented in this article.
@@ Acknowledgements.
R.M. would like to thank David Berenstein, Alejandra Castro, and Ioannis Papadimitriou for various discussions on related topics that, in part, inspired this project. R.M. would also like to thank Clifford Johnson and Jorge Zanelli for helpful discussions, as well as David Berenstein, Alejandra Castro, Adolfo Holguin, Sean McBride, Andrew Svesko, and Manus Visser for their comments on the manuscript. R.M. is grateful to the Centro de Estudios Científicos, Valdivia, and Pontificia Universidad Católica de Chile, Santiago, for their hospitality while part of this manuscript was being written. R.M.’s research was supported in part by the Department of Energy under grant DE-SC 0011702, and by the Chilean Fulbright Commission and ANID through the Beca Igualdad de Oportunidades ##\## 56180016.

### v3

@@ 5 Comments on cosmological constant as a thermodynamic variable
where $d$ is the dimension of the bulk spacetime. [^10] For reviews on extended black hole thermodynamics, see Mann4 ; Mann5 . Other relevant literature on the subject includes Mann2 ; Mann3 ; Clifford2 ; Kastor1 . Clearly, the notion of pressure $p_{\Lambda}$ does not correspond to the pressure of the dual field theory. For this reason, the physical interpretation of the Smarr formula was investigated from a holographic perspective in Karch .
@@ 6 Discussion
Our construction was restricted to the zeroth-order derivative expansion of the fluid variables, since our primary goal was to generalize the Euler equation. A natural direction for future work would be to extend this analysis to higher-derivative terms in the effective action, in order to investigate whether the introduction of the new variable modifies the transport coefficients.
@@
It’s well-known that the Schwarzian action captures the near-horizon dynamics of near-extremal black holes Amhed ; Stanford . In contrast, our effective action given in (42) is an off-shell formulation that captures not only the one-point functions (the stress tensor), but also the Ward identities (the Navier-Stokes equations without dissipation). In this way, following closely the work of Me , it should be possible to derive a decoupling limit from the action (70), which encodes the dynamics of a holographic charged fluid, and connect it to the one-dimensional anomalous perfect fluid of Jensen2 . In fact, the author of Me already achieved this decoupling limit for a subclass of solutions of a holographic CFT3, although using a different set of variables from the fluid variables.
@@
Similarly, the isometries of the degenerate target metric $G_{MN}$ given in (24) are closely related to Carrollian diffeomorphisms. Motivated by this, it should be possible to formulate an effective action for Carrollian fluids CR1 ; CR2 ; CR3 ; CR4 ; CR5 ; CR6 ; CR7 ; CR8 . In this context, Carrollian diffeomorphisms would serve both as gauge symmetries of the effective action and as spacetime symmetries of the fluid’s background. [^11] This perspective aligns naturally with the Gauge/Gravity duality Liu5 , as it likely follows from taking a flat limit of the fluid. There is also a body of work exploring the connection between Carrollian fluids and horizon symmetries CR9 ; CR10 ; CR11 ; CR12 ; CR13 . With this action principle in place, the Smarr formula for asymptotically flat black holes should follow in a manner similar to the derivation presented in this article.
@@ Acknowledgements.
R.M. would like to thank David Berenstein, Alejandra Castro, and Ioannis Papadimitriou for various discussions on related topics that, in part, inspired this project. R.M. would also like to thank Clifford Johnson and Jorge Zanelli for helpful discussions, as well as David Berenstein, Alejandra Castro, Adolfo Holguin, Sean McBride, Andrew Svesko, and Manus Visser for their comments on the manuscript. R.M. is grateful to the Centro de Estudios Científicos, Valdivia, and Pontificia Universidad Católica de Chile, Santiago, for their hospitality while part of this manuscript was being written. R.M.’s research was supported in part by the Department of Energy under grant DE-SC 0011702, and by the Chilean Fulbright Commission and ANID through the Beca Igualdad de Oportunidades #\# 56180016.